Item Account Debit Credit Accounts Payable Accounts Receivable Factory Overhead-Making Department Materials Work in Process-Making Department Work in Process-Packing Department

Entries for Process Cost System

Preston & Grover Soap Company manufactures powdered detergent. Phosphate is placed in process in the Making Department, where it is turned into granulars. The output of Making is transferred to the Packing Department, where packaging is added at the beginning of the process. On July 1, Preston & Grover Soap Company had the following inventories:

Finished Goods $27,550
Work in Process—Making 10,710
Work in Process—Packing 13,960
Materials 6,050

Departmental accounts are maintained for factory overhead, which both have zero balances on July 1.

Manufacturing operations for July are summarized as follows:

a. Materials purchased on account $343,170
b. Materials requisitioned for use
Phosphate—Making Department $226,690
Packaging—Packing Department 78,850
Indirect materials—Making Department 8,870
Indirect materials—Packing Department 3,180
c. Labor used
Direct labor—Making Department $161,950
Direct labor—Packing Department 109,310
Indirect labor—Making Department 31,360
Indirect labor—Packing Department 56,220
d. Depreciation charged on fixed assets
Making Department $29,570
Packing Department 24,420
e. Expired prepaid factory insurance
Making Department $5,600
Packing Department 2,240
f. Applied factory overhead
Making Department $77,280
Packing Department 85,390
g. Production costs transferred from Making Department to Packing Department $467,260
h. Production costs transferred from Packing Department to Finished Goods $735,390
i. Cost of goods sold during the period $738,080

Required:

1. Journalize the entries to record the operations, identifying each entry by letter. For a compound transaction, if an amount box does not require an entry, leave it blank.

Item Account Debit Credit
a.
b.
c.
d.
e.
f.
g.
h.
i.

2. Compute the July 31 balances of the inventory accounts.

Materials $
Work in Process—Making Department $
Work in Process—Packing Department $
Finished Goods $

3. Compute the July 31 balances of the factory overhead accounts. If required, use the minus sign to indicate a credit balance.

Factory Overhead—Making Department $
Factory Overhead—Packing Department $
Answer #1

1)

Date Account title Debit credit
a Raw material inventory 343170
Accounts payable 343170
b Work in process-making 226690
Work in process-Packing 78850
Factory overhead -Making 8870
Factory overhead -pacling 3180
Raw material inventory 317590
c Work in process-making 161950
Work in process-Packing 109310
Factory overhead -Making 31360
Factory overhead -pacling 56220
Factory labor 358840
d Factory overhead -making 29570
Factory overhead -packing 24420
Accumulated depreciation -fixed asset 53990
e Factory overhead -making 5600
Factory overhead -packing 2240
Prepaid insurance 7840
f work in process-making 77280
work in process-packing 85390
Factory overhead-making 77280
Factory overhead -packing 85390
g WOrk in process -packing 467260
work in process-making 467260
h Finished goods inventory 735390
work in process-packing 735390
I cost of goods sold 738,080
Finished goods inventory 738080

2)

Material =Beginning +purchase- ending

            = 6050+343170-317590

            = 31630

work in process- making =Beginning +current month cost-ending

                  = 10710+[226690+161950+77280]-467260

                 = 10710 +465920-467260

                = 9370

work in process -packing = 13960+ [78850+109310+85390]+467260-735390

           = 13960+ 273550+467260-735390

            = 19380

Finished goods = beginning + transfered from packing - cost of goods sold

          = 27550+735390-738080

          = 24860

3)Factory overhead = Actual -applied

     = [8870+31360+29570+5600]-77280

    = 75400-77280

        =- 1880

Facotry overhead -packing =[3180+56220+24420+2240]-85390

     = 86060-85390

    = 670